whitemonsterbunny17 (whitemonsterbunny17):

Mooooore contrapositives. .-.

whitemonsterbunny17 (whitemonsterbunny17):

If y is greater than 3, then y does not equal 4. I know that the contrapositive would start out as "If y equals 4," but I have no idea what to do with the greater than symbol thingy.

OpenStudy (ajspeller):

say it is not greater than 3

whitemonsterbunny17 (whitemonsterbunny17):

What would that look like? This? 

OpenStudy (ajspeller):

yes

OpenStudy (ajspeller):

or != or <>

OpenStudy (anonymous):

\[ \neg (x\gt y) \iff x\leq y \]

OpenStudy (anonymous):

\[ \neg (x\gt y) \iff x=y\lor x\lt y\iff x\leq y \]

whitemonsterbunny17 (whitemonsterbunny17):

All these symbols are confusing me. o-O

whitemonsterbunny17 (whitemonsterbunny17):

So just put a slash through the inequality symbol, right?

whitemonsterbunny17 (whitemonsterbunny17):

@ganeshie8

ganeshie8 (ganeshie8):

`If y is greater than 3, then y does not equal 4. ` your given statement is in words or symbols ?

ganeshie8 (ganeshie8):

statement : if \(y \gt 3\), then \(y \ne 4\) like this ?

whitemonsterbunny17 (whitemonsterbunny17):

Yes. ^.^

ganeshie8 (ganeshie8):

then you're absolutely right! just swap and negate

ganeshie8 (ganeshie8):

statement : if \(y \gt 3\), then \(y \ne 4\) contrapositive : if \(y = 4\), then \(y \not \gt 3\)

ganeshie8 (ganeshie8):

thats technically correct, but doenst that look ugly ?

whitemonsterbunny17 (whitemonsterbunny17):

Yeah, it looks strange, that's why I was confused. xD

ganeshie8 (ganeshie8):

we may replace \(\not \gt\) by the familiar and cute symbol \(\le \)

ganeshie8 (ganeshie8):

contrapositive : if \(y = 4\), then \(y \le 3\) see if this looks good

ganeshie8 (ganeshie8):

saying you are "not taller" than me is same as saying you're "below or equal to" my height

whitemonsterbunny17 (whitemonsterbunny17):

Ah, true. Thanks!! \(\Large\color{magenta}{\heartsuit}\)

ganeshie8 (ganeshie8):

np :)